What is the next three terms of the sequence, 49,44,39,34,29
Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

Check the difference between terms.

Answer by Theo(13342)

looks like you're subtracting 5 from each preceding term to get the current term starting with the first term being the preceding term.
49 - 5 = 44 - 5 = 39 - 5 = 35 - 5 = 29
the sequence formula to use if An = A1 + (n-1) * d
A1 is 49
d is -5
when n = 1, the formula becomes A1 = 49 + 0 * -5 = 49
when n = 2, the formula becomes A2 = 49 + 1 * -5 = 44
when n = 3, the formula becomes A3 = 49 + 2 * -5 = 39
etc.....